# A Deep Learning-Based Supervised Transfer Learning Framework for DOA Estimation with Array Imperfections

### v2

@@ V SIMULATIONS AND ANALYSES
As the DOA estimation system is a typical electronic system, the performance of the model in DOA estimation is influenced by a combination of multiple parameters. We summarize the primary factors as follows:
  1. AS1
    SNR and the number of snapshots. These two parameters are the most extensively discussed in DOA estimation fields. It is generally accepted that high SNR and a large snapshots lead to an ideal signal environment, thereby enhancing algorithm performance.
  2. AS2
    Number of array elements $\boldsymbol{M}$ and number of incident signals $\boldsymbol{K}$. A larger $M$ and a smaller $K$ facilitate more accurate DOA estimation.
  3. AS3
    Angular Separation and Range of incident signals. Discussions regarding the angular parameters of incident signals are relatively limited; however, these factors significantly impact a model’s performance. In general, smaller angular separations result in more similar steering vectors, leading to increasingly difficult DOA estimation. This situation poses a considerable challenge to the model’s angular resolution. Additionally, the range of incoming signals can affect the DOA estimation performance of specific array types (such as $\theta$ for ULA or $\phi$ for UCA); greater angles of signal incidence lead to reduced angular resolution for these arrays.
In this section, multiple distinct simulation scenarios were constructed based on the settings of the aforementioned parameters. We demonstrate the effectiveness of the proposed model and transfer learning approach through extensive simulation experiments. We begin with a detailed description of the experimental setup in Subsection V-A. Subsequently, the performance of our proposed model under these setups is reported in Subsection V-B. Furthermore, we demonstrate that our proposed model exhibits significant performance degradation in scenarios involving array imperfections, which motivates the introduction of transfer learning techniques to address the challenges posed by array imperfections. The simulation results are presented in Subsection V-C.
A wide range of classic and state-of-the-art algorithms were selected to demonstrate the effectiveness of the proposed model. We first selected several classic subspace-based approaches as baseline methods for comparison, such as MUSIC [17] and Unity-ESPRIT algorithms [16]. Compressive sensing algorithm $\ell_{1}$-SVD algorithm [24, 63] is considerd, which has strong performance in low SNR and small snapshots scenarios. In addition, a series of deep learning methods and their variants that integrate traditional algorithms were used for comparative analysis. We selected spatial spectrum estimation CNN algorithm (SPE-CNN) [1], SubspaceNet [35], ASL-2 [33], and Learning-SPICE [34] for comparison. SPE-CNN is recognized as a well-established deep learning-based method and SubspaceNet can be understood as a subspace algorithm implemented using deep learning networks. Furthermore, ASL-2 and Learning-SPICE, which representing the latest algorithms that combine deep learning and compressive sensing, are expected to perform well in small snapshots scenarios.
@@ V-A Experimental Setup
1) Simulation Scenarios: Throughout this study, we consider four distinct scenarios to evaluate the performance of our proposed algorithm:
  • Scenario 1: A ULA with $M=8$ and $K=3$.
  • Scenario 2: A ULA with $M=8$ and $K=7$.
  • Scenario 3: A ULA with $M=16$ and $K=3$.
  • Scenario 4: A UCA with $M=12$ and $K=5$.
@@
Figure 4: DOA estimation under Scenarios 1 with $SNR=-5db$ and $snap=10$. Three random incident signals are generate and we focus on the first one. The points represent the estimated DOA provided by the algorithms while the circles denote the true DOAs.
@@
TABLE I: CDF-Based Statistical Evaluation results under Scenarios 1 with $snap=10$ and $SNR=-5$.
Algorithms Ours MUSIC Unity-ESPRIT $\ell_{1}$-SVD SPE-CNN ASL-2 SubspaceNet Learning-SPICE
$\text{CDF}(3)$ 55.36% 54.10% 48.90% 44.30% 60.11% 33.49% 52.63% 68.08%
$\text{CDF}(10)$ 89.20% 73.43% 80.31% 82.93% 81.35% 55.68% 87.35% 87.51%
CDF10% $0.455^{\circ}$ $0.410^{\circ}$ $0.526^{\circ}$ $0.638^{\circ}$ $0.368^{\circ}$ $0.766^{\circ}$ $0.501^{\circ}$ $\mathbf{0.309^{\circ}}$
CDF90% $\mathbf{10.63^{\circ}}$ $38.13^{\circ}$ $24.22^{\circ}$ $20.91^{\circ}$ $29.66^{\circ}$ $55.20^{\circ}$ $13.55^{\circ}$ $16.94^{\circ}$
Scenario 1 represents a standard multi-target DOA estimation scenario, which is widely adopted for evaluating model performance. Subsequently, to further investigate whether the model’s performance deteriorates when handling a larger number of incident angles, Scenario 2 is introduced.
Scenario 3 represents a scenario with more array elements, which aims to achieve precise DOA estimation. In such cases, algorithms typically demand significantly greater computational resources. Thus, maintaining accuracy while minimizing computational overhead becomes a critical challenge in this scenario.
@@
Scenario 4 involves the use of a UCA for two-dimensional localization. This scenario evaluates the model’s capability to extend from one-dimensional to two-dimensional tasks, and tests its performance in 2D DOA estimation.
The aforementioned scenarios essentially modify the parameters of AS2. Within each scenario, data with varying SNR, snapshot numbers, angular separations and angular ranges are thoroughly tested. This comprehensive evaluation aims to provide a thorough assessment of the model’s performance, including model’s adaptability to low SNR and small snapshots scenarios, model’s angular resolution capabilities, and model’s angular estimation range in various scenarios.
2) Hyperparameter Settings: In the developed DOA-ViT model, the dimensionality of the embedded vectors $D$ is fixed at 768, while the dimension of each head $d_{k}$ is set to 64. Additionally, our model consists of 6 Transformer encoders and the MHSA module is configured with 12 attention heads. Within each encoder, the expansion factor of the MLP layers is set to 4.
@@
Under each configuration, 50,000 samples are generated for model training, and 20,000 samples are generated for model validation. For a ULA, the incident angles $\theta$ are randomly generated within the range $[-60^{\circ},60^{\circ}]$ in the training and validation sets. And for a UCA, the incident angles $\theta$ are randomly generated within $[-180^{\circ},180^{\circ}]$, and $\phi$ values are randomly generated within $[0^{\circ},60^{\circ}]$.
The Adam optimizer is configured with a learning rate of $0.0001$; betas values $(0.9,0.999)$ are used for the first and second moment estimates, which is considered effective in most scenarios. The number of training epochs is set to 500. However, if the validation loss does not improve over the course of 30 epochs, early stopping is triggered to halt the training.
@@
For the sake of simplicity and fairness, we train each deep learning model for comparison using the same dataset and optimizer to ensure consistency in the experiments and comparability of the results. For more detailed implementation and configuration specifics, the complete code is available in our open-source repository [^1] https://github.com/zzb-nice/DOA_est_Master, providing further insight into the structure and parameter settings employed.
V-B DOA estimation performance based on ViT model
@@ V-B1 Simulation results in Scenario 1
A series of simulation experiments were conducted to validate the effectiveness of the proposed ViT model. First, we evaluate the performance of the model in Scenario 1.
Visual Performance Demonstration: In order to provide a visual comparison of the performance of various algorithms in Scenario 1, specifically in the low SNR and small snapshot conditions, we visually demonstrate the DOA estimation results of multiple algorithms in Fig 4. Scenario 1 with SNR set to -5 dB and snapshots number set to 10 is considered, which serves as the low SNR, small snapshot scenario. The absolute angular error of the model’s DOA estimates is plotted on the right side. Through a direct comparison, the proposed method demonstrates superior performance, with the majority of estimation errors being smaller than $25^{\circ}$ in such a low SNR, small snapshot scenario. In contrast, all other algorithms exhibit severe deviations with absolute errors exceeding 50 degrees.
Figure 5: RMSE $vs.$ SNR results under Scenarios 1 with $snap=10$. The incident signal’s DOA values are randomly generated with the minimum angular interval set to 3 degrees.
RMSE $\boldsymbol{vs.}$ SNR: To effectively evaluate the model’s performance, it is essential to assess its statistical performance. We demonstrate the Root Mean Squared Error (RMSE) of estimated DOA values under varying SNR and snapshots (AS1).
In Fig. 5, we investigate the RMSE vs. SNR results based on 5,000 simulations. It can be seen that the proposed method achieves the best performance when the incident angles are randomly generated, which can be regarded as the most general case. Learning-SPICE closely approaches the performance of the proposed model as SNR increase ($SNR=0db$ or $SNR=5db$). Unity-ESPRIT and SPE-CNN perform relatively well at high SNRs, while SubspaceNet achieves better results in low SNR scenarios. $\ell_{1}$-SVD exhibits consistently stable performance, and MUSIC and ASL-2 exhibit inferior performance compared with the aforementioned algorithms.
Furthermore, in order to provide a more intuitive demonstration of the model’s statistical performance and a more comprehensive evaluation of its overall performance, the cumulative distribution function (CDF) of the DOA estimation absolute error is illustrated to assess the model’s performance, which describes the probability distribution of the absolute error. As shown in Fig. 6, 5,000 simulations are conducted in each condition, where the incident signal’s DOA values are randomly generated with the minimum angular interval set to 3 degrees.
Through the CDF, it can be observed that under higher SNR conditions (e.g., $SNR=0db$ or $5db$), the proposed model, Unity-ESPRIT, SPE-CNN and Learning-SPICE exhibited commendable performance. However, compared with other algorithms, the proposed model exhibits a lower cumulative probability of small absolute errors, and it shows a distinct advantage in larger absolute error ranges. This trend is more evident in low SNR scenarios like $SNR\leq 0db$, as detailed in the following analysis.
@@
Figure 6: CDF of the absolute error under Scenarios 1 with $snap=10$ and various $SNR$, for (a) $SNR=0db$, (b) $SNR=5db$, (c) $SNR=-5db$ and (d) $SNR=-10db$.
@@
A representative scenario is depicted in Fig. 6, where it can be observed that the distribution of absolute errors within the range of $0^{\circ}$ to $3^{\circ}$ for the proposed model is slightly lower than those of Learning-SPICE, SPE-CNN and MUSIC, while exhibiting advantage over the SubspaceNet, Unity-ESPRIT, $\ell_{1}$-SVD and ASL-2 algorithms. However, the proposed model’s probability of absolute errors exceeding $10^{\circ}$ is substantially lower, which indicates a reduced frequency of high-error occurrences, highlighting our model’s superior robustness relative to other approaches.
Figure 7: RMSE $vs.$ SNR results under Scenarios 1 with $snap=10$. The dashed line in the figure represents Scene 1, while the solid line represents Scene 2. In Scene 1, the true DOA values are [10°, 13°, 16°], and in Scene 2, the true DOA values are [10°, 20°, 30°].
Figure 8: RMSE $vs.$ SNR results under Scenarios 2 with $snap=10$. The incident signals are randomly generated with the minimum angular interval set to 3 degrees.
Figure 9: RMSE $vs.$ snap results under Scenarios 1 with $SNR=-10db$. The dashed line in the figure represents Scene 1, while the solid line represents Scene 2. In Scene 1, the true DOA values are [5.1°, 30.3°, 54.5°], and in Scene 2, the true DOA values are [-9.9°, 15.3°, 39.5°].
Figure 10: RMSE $vs.$ SNR results under Scenarios 3 with $snap=10$. The incident signals are randomly generated with the minimum angular interval set to 3 degrees.
Furthermore, through the CDF, a series of statistical measures related to the absolute error can be introduced. For example, CDF50% represents the absolute error corresponding to a cumulative probability of ${50\%}$. As a quantile-based measure of absolute error, it differs from the commonly used mean absolute error (MSE), which represents the mean of absolute errors. Both serve as statistical measures of absolute error and hold statistical significance. Similarly, CDF(5) refers to the probability that the absolute error is less than 5, which can also reflect the model’s performance.
The results of the CDF-based statistical evaluation for various models are presented in Table I. From the table, it can be observed that Learning-SPICE demonstrates superior performance at $\text{CDF}(3)$ and CDF10%, achieving the highest DOA estimation accuracy, where the algorithm’s estimated values are most likely to be close to the true values. However, our proposed model demonstrates the best performance at $\text{CDF}(10)$ and CDF90%, showing greater robustness by reducing the occurrence of large errors. Due to the fact that RMSE penalizes larger estimation errors severely, the proposed model demonstrates superior performance in terms of RMSE compared with the Learning-SPICE algorithm.
To further demonstrate the performance of the proposed model, we selected four representative models mentioned above for testing in smaller angular interval scenarios. We consider two Scenes with different angular separations (AS3): Scene 1 represents the scenario with signals incident at [10°, 13°, 16°], while Scene 2 represents the scenario with signals incident at [10°, 20°, 30°]. In Scene 1, the angular interval between incident signals is smaller, which imposes higher demands on the model’s angular super-resolution capability.
In general, since smaller angular intervals result in more similar steer vectors, algorithms based on linear space and compress sensing will perform better in Scene 2 than in Scene 1. On the other hand, deep learning algorithms are fundamentally designed to minimize the loss function. By minimizing the loss function across the entire training set, deep learning algorithms may achieve improved performance at smaller angular intervals.
[…]
RMSE $\boldsymbol{vs.}$ snapshots: To further demonstrate the model’s performance under varying snapshots (AS1) and angular ranges (AS3) for a ULA, we conducted simulation experiments in different snapshot scenarios. Similar to the previous experiment, we empirically evaluate the RMSE by averaging over 5,000 Monte Carlo trials. Two distinct scenarios are considered to investigate the model’s performance under different angular ranges. Scene 1 represents the scenario with signals incident at [5.1°, 30.3°, 54.5°], while Scene 2 represents the scenario with signals incident at [-9.9°, 15.3°, 39.5°]. The demands on the model’s angular super-resolution are alleviated and the angular intervals are kept the same in two scenes. Thus, the larger angular range in Scene 2 leads to a lower angular resolution, which has a negative impact on the performance of the algorithms.
As shown in Fig. 10, a clear trend can be observed: for any given algorithm, the performance in Scene 1 is inferior to that in Scene 2, which can be attributed to the more challenging conditions in Scene 1. Furthermore, the proposed DOA-ViT model generally outperforms the other algorithms in this low SNR scenario. In Scene 1, our model achieves the best performance, surpassing the results achieved by MUSIC, $\ell_{1}$-SVD, Learning-SPICE, and Unity-ESPRIT. In Scene 2, the proposed model also attains the best performance, slightly outperforming the Learning-SPICE algorithm. These results highlight the superior effectiveness of our approach.

### v5

@@ V SIMULATIONS AND ANALYSES
In this section, we evaluate the performance of the proposed scheme through extensive simulations, comprehensively validating its effectiveness and robustness under varying critical factors, specifically SNR, number of snapshots $T$, number of array elements $M$, number of incident signals $K$, and angular separation.
Comprehensive comparisons are conducted with a wide range of classic and state-of-the-art model-driven methods, including high-resolution subspace methods such as MUSIC and Unity-ESPRIT algorithms, as well as the compressive sensing-based $\ell_{1}$-SVD approach [18, 45], which has strong performance in low SNR and small snapshots scenarios. Additionally, a series of deep learning methods and their variants that integrate traditional algorithms are considered, including the spatial spectrum estimation CNN algorithm (SPE-CNN) [34], SubspaceNet [39], ASL-2 [37], and Learning-SPICE [38]. SPE-CNN is recognized as a well-established deep learning-based method and SubspaceNet can be understood as a subspace algorithm implemented using deep learning networks. Furthermore, ASL-2 and Learning-SPICE, which representing the latest algorithms that combine deep learning and compressive sensing, are expected to perform well in limited snapshots scenarios.
We begin with a detailed description of the experimental setup in Subsection V-A. The performance of the proposed model under these setups is then presented in Subsection V-B. Furthermore, in Subsection V-C, we validate the effectiveness of transfer learning strategies through simulations conducted in practical scenarios involving array imperfections.
@@ V-A Experimental Setup
TABLE I: Simulation Scenarios and Parameter Configurations
Scen. Setting $(M,K)$ Configuration & Test Conditions
1 ULA $(8,3)$ DOA: Uniform, Equidistant
Params [1]: SNR =5=-5 dB, $T=10$
Params [2]: SNR $\in[-20,5]$ dB, $T=10$
Params [3]: $T\in[10,100]$, SNR=10=-10 dB
2 ULA $(8,7)$ DOA: Uniform
Params [1]: SNR $\in[-20,5]$ dB, $T=10$
3 ULA $(16,3)$ DOA: Uniform
Params [1]: SNR $\in[-20,5]$ dB, $T=10$
4 UCA $(12,5)$ DOA: Deterministic, Uniform
Params [1]: SNR =5=-5 dB, $T=50$
Params [2]: SNR =5=5 dB, $T=10$
  • Uniform: DOAs are sampled from a uniform distribution, subject to a minimum angular separation of $3^{\circ}$ to avoid source clustering.
  • Equidistant: DOAs are equally spaced across the entire angular sector.
  • Deterministic: A predetermined set of specific DOAs.
1) Simulation Scenarios: To facilitate a holistic assessment of the algorithm’s performance, we conduct simulations across four distinct scenarios, as summarized in Table I.
@@ 4 UCA $(12,5)$ DOA: Deterministic, Uniform
Figure 5: DOA estimation results under Scenarios 1 with SNR=5dbSNR=-5db and $snap=10$. Three incident signals are generate and we focus on the first one. The points represent the estimated DOA while the circles denote the true DOAs.
@@
TABLE II: CDF-Based Statistical Evaluation results under Scenarios 1 with SNR=5SNR=-5 and $snap=10$.
Method Miss Prob. $\downarrow$ OSPA Error RMSE () MAE () Acc. (%) ECDF Quantile (Raw) ECDF Quantile (Matched)
Linear $\downarrow$ Square $\downarrow$ Raw $\downarrow$ Matched $\downarrow$ Raw $\downarrow$ Matched $\downarrow$ Raw $\uparrow$ Matched $\uparrow$ 10% $\downarrow$ 90% $\downarrow$ 10% $\downarrow$ 90% $\downarrow$
MUSIC 0.001 6.65 9.52 20.39 12.11 10.81 6.68 74.3 80.9 0.43 37.31 0.38 30.00
Unity-ESPRIT 0.000 6.27 8.46 17.09 10.52 8.92 6.29 78.8 82.8 0.50 26.62 0.43 20.82
$\ell_{1}$-SVD 0.088 6.35 8.36 14.33 9.36 7.60 5.63 80.8 85.4 0.62 20.04 0.58 12.38
SPE-CNN 0.000 5.53 8.08 18.79 10.54 8.84 5.55 80.6 84.9 0.39 29.47 0.32 25.99
Learning-SPICE 0.000 4.47 6.34 15.09 9.07 6.43 4.50 87.4 89.1 0.31 18.98 0.28 13.82
ASL-2 0.049 11.62 14.28 31.66 15.57 19.91 11.06 54.6 60.7 0.77 60.00 0.62 30.00
SubspaceNet 0.000 5.15 6.79 12.71 8.88 6.47 5.16 85.8 88.3 0.52 16.30 0.48 12.27
TransDOA (ours) 0.000 4.68 5.99 9.65 7.73 5.18 4.69 88.2 89.0 0.46 11.60 0.45 10.93
Scenario 1 establishes a standard benchmark for multi-source estimation for a ULA. Scenario 2 considers a high-density source environment where the number of signals is near the degrees of freedom of the array.
@@
Scenario 3 represents a scenario with more array elements, which aims to achieve precise DOA estimation. Scenario 4 involves the use of a UCA for two-dimensional localization. This scenario evaluates the model’s capability to extend from one-dimensional to two-dimensional tasks.
To ensure a comprehensive evaluation, we performed simulations across a range of angular configurations, SNRs, and snapshots in each scenario. The specific parameter settings are detailed in Table I.
2) Implementation Details: The proposed TransDOA model comprises 6 Transformer encoder layers. the embedding dimension $D$ is set to 768 and each MHSA module employs 12 attention heads. Within each encoder, the expansion factor of the feed-forward network is set to 4.
@@
Under each simulation scenario, a total of 50,000 training samples and 20,000 validation samples are generated. For the ULA, the DOAs $\theta$ are randomly generated within the range $60^{\circ},60^{\circ}]$ in the training and validation sets. And for the UCA, the azimuth angle $\theta$ and elevation angle $\phi$ are randomly generated within $180^{\circ},180^{\circ}]$ and $0^{\circ},60^{\circ}]$, respectively.
Model optimization is performed using the Adam optimizer with a learning rate of $10^{-4}$ and exponential decay rates $(\beta_{1},\beta_{2})=(0.9,0.999)$. The number of training epochs is set to 500, with an early stopping mechanism triggered if the validation loss fails to improve for 30 epochs.
@@
To ensure a fair and rigorous comparison, all deep learning models are trained using the identical dataset and optimization protocol, thereby maintaining experimental consistency and the comparability of results. For more detailed implementation and configuration specifics, the complete code is available in our open-source repository [^1] https://github.com/zzb-nice/DOA_est_Master, providing further insight into the structure and parameter settings employed.
V-B DOA estimation performance based on TransDOA model
@@ V-B1 Simulation results in Scenario 1
To validate the effectiveness of the proposed TransDOA model, we first evaluate its performance in Scenario 1 through a series of targeted experiments.
Visual Performance Demonstration: To visually evaluate the robustness of various algorithms under challenging signal environments, Fig. 5 illustrates the DOA estimation results in Scenario 1 with SNR= -5 dB and TT=10, which serves as the low SNR, limited snapshot scenario. The right-hand plot in the figure depicts the absolute angular error of each method. As illustrated, the proposed method demonstrates superior performance, with most estimation errors remaining below $25^{\circ}$, whereas all other algorithms suffer from severe deviations, with absolute errors exceeding 50 degrees.
Figure 6: RMSE $vs.$ SNR results under Scenarios 1 with Uniform configuration and Params [2] specified in Table I.
RMSE $\boldsymbol{vs.}$ SNR: To evaluate the statistical performance of the model, we report the Root Mean Squared Error (RMSE) of the estimated DOAs under varying SNR conditions based on 5,000 simulations. As illustrated in Fig. 6, The proposed method achieves the best RMSE performance when DOAs are randomly generated (Uniform configuration), which can be regarded as the most general case. Learning-SPICE closely approaches the performance of the proposed model as SNR increase ($SNR=0db$ or $SNR=5db$). Unity-ESPRIT and SPE-CNN perform relatively well at high SNRs, while SubspaceNet achieves better results in low SNR scenarios. $\ell_{1}$-SVD exhibits consistently stable performance, and MUSIC and ASL-2 exhibit inferior performance compared with the aforementioned algorithms.
To provide a more granular evaluation of the statistical characteristics, the empirical cumulative distribution function (ECDF) of the absolute DOA estimation error is illustrated to assess the model’s performance. As shown in Fig. 7, The ECDF provides insight into the error distribution across 5,000 simulations under Scenarios 1 with Uniform configuration.
As observed from the ECDF curves, while the proposed model exhibits a slightly lower cumulative probability within the high-precision regime, it demonstrates a distinct advantage in larger error ranges. This indicates that the proposed method prioritizes global robustness, making it significantly less prone to catastrophic estimation failures. This trend is particularly evident in low SNR scenarios (e.g., $SNR\leq 0db$).
@@
Figure 7: ECDF of the absolute error under Scenarios 1 with with Uniform configuration and Params [2], for (a) $SNR=0db$, (b) $SNR=5db$, (c) SNR=5dbSNR=-5db and (d) SNR=10dbSNR=-10db.
@@
A representative scenario is depicted in Fig. 7. within the high-precision regime ($0^{\circ}-3^{\circ}$), the cumulative probability of the proposed model is slightly lower than that of Learning-SPICE, SPE-CNN and MUSIC, while exhibiting advantage over the SubspaceNet, Unity-ESPRIT, and other algorithms. However, the probability of absolute errors exceeding $10^{\circ}$ is substantially reduced. This lower frequency of high-error occurrences highlights the model’s superior reliability and its ability to suppress outliers in challenging environments.
Figure 8: RMSE $vs.$ SNR results under Scenarios 2 with Uniform configuration and Params [1].
Figure 9: RMSE $vs.$ SNR results under Scenarios 3 with Uniform configuration and Params [1].
To rigorously evaluate the model’s performance, we move beyond traditional sorting-based pairing schemes, which fail to accommodate cardinality mismatches and introduce systematic bias by discarding samples with missed detections. Instead, we utilize the Hungarian algorithm for optimal source association, with estimation error capped at $30^{\circ}$ to prevent extreme outliers from distorting the statistical results. Evaluation is conducted across both ’Raw’ and ’Matched’ configurations using a comprehensive suite of metrics: OSPA [46] (to jointly assess cardinality and localization accuracy), MAE, RMSE, Accuracy (Acc., success ratio within a $10^{\circ}$ tolerance), and the 10th/90th percentiles of the ECDF. These quantiles respectively characterize the high-precision capability and the robustness boundary of the model.
The performance results for Scenario 1 (Uniform configuration, Params [1]) are detailed in Table II. In the table, algorithms with non-zero miss probability are shaded in gray, signifying a potential risk of target loss. Directional arrows ($\downarrow$ / $\uparrow$) are employed to indicate whether a higher or lower value represents superior performance. For clarity, the optimal values for each metric are bolded, while the second-best results are underlined. It is observed that the proposed algorithm consistently achieves the best or second-best performance across the majority of metrics, demonstrating its competitive edge in both estimation accuracy and reliability.